Price per CD 0 10 20 70 80 90 30 40 50 60 Millions of CDs per Month

find tax incidence so CTI AND PTI
Answer #1

Answer
the tax shifts the supply curve from S1 to S2 so the vertical distance is the tax per unit
tax per unit =0.7-0.5=0.2
After tax equilibrium is at D=S2
where
consumer pay =$0.7 and Q=40
Producer pay =consumer pay -tax=0.7-0.2=0.5
before tax equilibrium is at D=S1
where
Q=50 and P=0.6
=====
CTI =consumer tax incidance =after tax consumer pay - before tax consumer pay =0.7-0.6=0.1

PTI=per unit tax - CTI=0.2-0.1=0.1
the CTI and PTI is $0.1 each